OpenStudy (anonymous):

It would take Nina 5 hours to move all the demolished bricks into the dumpster by herself. Together, Nina and Otis can move all of them in 3 hours. How fast could Otis move all the bricks by himself to the nearest tenth of an hour? a. 0.1 of an hour b. 0.2 of an hour c. 7.5 hours d. 8.0 hours

OpenStudy (anonymous):

@perl

OpenStudy (mayankdevnani):

i think option C is correct !!

OpenStudy (mayankdevnani):

@justsmile531 am i correct here ?? i mean can you check answer of this question

OpenStudy (anonymous):

I will let you know when I finish my quiz . I still have more questions to answer ; ughh I hate Algebra

OpenStudy (mayankdevnani):

please tag me if you know the answer,then i can explain you how to solve this question

OpenStudy (mayankdevnani):

okk.. @justsmile531

OpenStudy (anonymous):

You were correct

OpenStudy (mayankdevnani):

thanks !!!

OpenStudy (mayankdevnani):

Solution :- \[\large \bf \frac{1}{3}=\frac{1}{5}+\frac{1}{x}\] after simplify,you get x=15/2=7.5